# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 5.3 Agents for experiments planning
Building on the capabilities of ChemCrow and Coscientist in automating chemistry-related tasks, recent advances have focused on bridging the gap between virtual agents and physical laboratory environments. For example, Context-Aware Language Models for Science (CALMS)433, BioPlanner435, and CRISPR-GPT424 focus on giving support to researchers with wet-lab experimental design and data analysis.
@@
CALMS433 focuses on improving laboratory efficiency through the operation of instruments and management of complex experiments, employing conversational LLMs to interact with scientists during experiments. In addition, this agent can perform actions using lab equipment after lab equipment APIs have been provided to the agent as tools. CALMS was designed to enhance instrument usability and speed up scientific discovery, providing on-the-spot assistance for complex experimental setups, such as tomography scans, and enabling fully automated experiments. For instance, its capability was showcased through the operation of a real-world diffractometer. Although CALMS excelled in several tasks, a comparison between GPT-3.5 and Vicuna 1.5 revealed Vicuna’s limitations in handling tools.
@@
In contrast, BioPlanner435 significantly improves the efficiency of scientific experimentation by creating pseudocode representations of experimental procedures, showcasing AI’s capacity to streamline scientific workflows. Therefore, Rather than interacting directly with lab equipment through APIs, BioPlanner creates innovative experimental protocols that can be expanded upon within a laboratory setting. The initial step in BioPlanner’s process involves assessing the capability of LLMs to produce structured pseudocode based on detailed natural language descriptions of experimental procedures. In testing, BioPlanner successfully generated correct pseudocode for 59 out of 100 procedures using GPT-4, although the most common errors involved omitted units. Afterward, the authors used BioPlanned to generate a procedure for culturing an E.coli bacteria colony and storing it with cryopreservation, which ran successfully.
@@
Focusing on gene editing experiments, CRISPR-GPT424 is an agent developed to design experiments iteratively with constant human feedback. CRISPR-GPT424 aims to bridge the gap for non-experts by simplifying this process into manageable steps solvable by an LLM with access to useful tools. This agent operates in three modes based on user prompts: “Meta mode” provides predefined pipelines for common gene-editing scenarios; “Auto mode” uses the LLM to plan a sequence of tasks; and “Q&A mode” answers general questions about the experimental design. The authors demonstrate that based on human evaluations, CRISPR-GPT outperforms GPT-3.5 and GPT-4 in accuracy, reasoning, completeness, and conciseness. Additionally, they applied CRISPR-GPT to design real-world experiments for knocking out TGFBR1, SNAI1, BAX, and BCL2L1 in the human A375 cell line, achieving an editing efficiency of approximately 70% for each gene.
@@
Following the ideas of developing agents for automating experimental protocol generation, Ruan et al. 462 created a multi-agents system composed of 6 agents: Literature Scouter, Experiment Designer, Hardware Executor, Spectrum Analyzer, Separation Instructor, and Result Interpreter. The Large Language Models-based Reaction Development Framework (LLM-RDF)462 automates every step of the synthesis workflow. While other studies focus on the literature review463, 464, 465, HTS458, and reaction optimization466, 46, LLM-RDF can support researchers from literature search until the product purification. Using this system, the authors showed they could design a copper/TEMPO catalyzed alcohol oxidation reaction, optimize reaction conditions, engineer a scale-up, and purify the products, obtaining a yield of 86% and a purity >98% while producing 1 gram of product.
@@
Interestingly, despite acting on different fields and presenting different goals (while CALMS433 and LLM-RDF462 can autonomously use laboratory equipment, BioPlanner435 and CRISPR-GPT424 propose a protocol that needs to be carried out by a human), all studies implemented a “human-in-the-loop” approach. Maintaining the researcher as a vital piece of the development is a way to enhance reliability and mitigate agent limitations, such as errors and hallucinations. This approach also addresses risk and dual usage problems once a human can decide whether the agent’s suggestions are safe. 350, 467 On a slightly different direction, Organa346 fully automates the laboratory workload while providing feedback to the researcher and producing reports with the results, as discussed on Section 3.7.1.
@@ 5.4 Agents for automating cheminformatics tasks
Cheminformatics consists of applying information technology techniques to convert physicochemical information into knowledge. The process of solving cheminformatics problems commonly involves retrieving, processing, and analyzing chemical data.468 Getting inspiration from ChemCrow469 ideas, Chemistry Agent Connecting Tool Usage to Science (CACTUS)470 focused on assisting scientists by automating cheminformatics tasks. CACTUS automates the applications of multiple cheminformatics tools, such as property prediction and calculation, while maintaining the human-in-the-loop for molecular discovery. The authors investigated the performance of a diverse set of open-source LLMs, where Gemma-7B and Mistral-7B demonstrated superior performance against LLaMA-7B and Falcon-7B. In addition, the authors reported that adding domain-specific information in the prompt to align the agent to chemistry problems considerably increases model’s performance. For instance, predicting drug-likeness with a Gemma-7B agent improves the accuracy of $\sim$ 60% when aligning the agent. Prompt alignment improved the prediction of all properties they studied.
@@
Further illustrating the versatility of AI in scientific research and domain-specific tools usage is ChatMOF,437 which focuses on the prediction and generation of Metal-Organic Frameworks (MOFs). ChatMOF integrates MOF databases with its MOFTransformer471 predictor module, thereby showcasing the innovative use of genetic algorithms in guiding generative tasks from associated predictions. The authors showed that ChatMOF achieved an accuracy of $\sim$ 90% in search and prediction tasks while generative tasks have an accuracy of $\sim$ 70%. The genetic algorithm used by ChatMOF allows for the generation of a diverse array of MOF structures, which can be further refined based on specific properties requested by users. For instance, when prompted to, “generate structures with the largest surface area”, the system initially generated a broad distribution of structures with surface area centered in 3784 m2/g, and the GA evolves it to a narrower distribution with a peak at 5554 m2/g after only three generations. It is important to note that even though ChatMOF has access to a dataset of experimental values for MOFs, language model predictions guide their GA, and no further validation has been made. Lastly, Ansari and Moosavi 432 developed Eunomia, another domain-specific autonomous AI agent that leverages existing knowledge to answer questions about materials. Eunomia432 can use chemistry tools to access a variety of datasets, scientific papers and unstructured texts to extract and reason about material science information. The authors implemented a CoVe364 (Consistency Verification) scheme to evaluate the model’s answer and minimize hallucination. The authors showed that including CoVe increased the model’s precision by $\sim$ 20% when compared to previous methods such as an agent using ReAct only.369
@@
Promoting molecular discovery is a topic with great attention in the literature devoted to it and, as described extensively above, LLMs have leveraged a large amount of unstructured data to accelerate that discovery. Janakarajan et al. 436 discuss the advantages of using LLMs in fields such as de novo drug design, reaction chemistry, and property prediction, but they augment the LLM in IBM ChemChat, a chatbot with the capability of using common APIs and python packages commonly used daily by a cheminformatics researcher to access molecular information. ChemChat has access to tools such as Generative Toolkit for Scientific Discovery (GT4SD)472, a package with dozens of trained models generative models for science, rxn4chemistry473, a package for computing chemistry reactions tasks, HuggingMolecules474, a package developed to aggregate molecular property prediction LMs, and RDKit475, a package to manipulating molecules. Since ChemChat implements an agent in a chat-like environment, users can interactively refine design ideas. Despite being developed to target de novo drug design, ChemChat nonetheless is a multi-purpose platform that can be more broadly used for molecular discovery.
@@
In addition to the capabilities described above, LLM-based agents can empower users to tackle tasks that typically require extensive technical knowledge. In previous work, Wellawatte and Schwaller 286 and Gandhi and White 240 showed that including natural language explanations (NLE) in explainable AI (XAI) analysis can improve user understanding. More recently, Wellawatte and Schwaller 286 developed XpertAI286 to seamlessly integrate XAI techniques with LLMs to interpret and explain raw chemical data autonomously. Applying XAI techniques is usually restricted to technical experts but by integrating such techniques with an LLM-based agent to automate the workflow, the authors made XAI accessible to a wider audience.
@@
Their system receives raw data with labels for physicochemical properties. The raw data is used to compute human-interpretable descriptors and then calculate SHAP values or Z-scores for Local Interpretable Model-agnostic Explanations (LIME). By calculating SHAP values (or SHapley Addictive exPlanations) a value can be assigned to each feature, indicating its contribution to a model’s output. LIME interprets a model by making a local approximation, around a particular prediction, to indicate what factors have contributed to that prediction in the model. It may use, for example, a surrogate local linear regression fit to recognized features.240 In addition to XAI tools, XpertAI can search and leverage scientific literature to provide accessible NLEs. While ChatGPT provides scientific justifications with similar accuracy, its explanation is often too broad. On the other hand, XpertAI provides data-specific explanations and visual XAI plots to support its explanations.286
@@ 6 Challenges and opportunities
Quality and availability of data are critical factors that influence the efficacy of LLMs. Indeed, scaling both the model size and the amount of training data used has proven to improve capabilities.476 However, current AI models are not trained on large amounts of chemical data, which limits their capabilities to reason about advanced chemical concepts.137
There are two types of datasets commonly used to train LLMs: unlabeled and labeled datasets. Unlabeled datasets, or pretraining data, are used during the semi-supervised training, which focuses on creating a “prior belief” about a molecule. Currently, we have huge datasets composed of hypothetical and/or theoretical data. When a model is trained on data that is not grounded in real chemical information, this might cause the model to learn a wrong prior belief.477
Labeled datasets, often used in benchmarks, also suffer from their inclusion of hypothetical and calculated data. Benchmarks are necessary for quantifying improvements in AI modeling and prediction within a competitive field. However, dominant benchmarks like MoleculeNet,55 have significant limitations that may restrict the generalizability and applicability of evolving models. In his blog, Walters 142 brings to light numerous errors and inconsistencies within the MoleculeNet data, which substantially impact model performance and reliability.138, 139, 140, 141. Walters also argues that the properties present in these benchmarks do not directly correlate with real chemistry improvement. As such, new benchmarks need to translate to practical chemistry problems directly. For instance, increasing accuracy in predicting LogP is not necessarily mapped to drugs with greater bioavailability. Some promising work has come from the Therapeutic Common Data (TDC)145, 144 includes data from actual therapeutic essays, providing a more practical foundation for model training.
The community continues to work to organize and curate datasets to make data ready for LLMs training and evaluation. Scientific benchmarks163, 430, 478, repositories with curated datasets136, and packages for model evaluation137 have been developed. However, the challenges concerning grounded truth and consistent datasets remain. With advancements in scientific document processing,316 there is now the opportunity to obtain new datasets from peer-reviewed scientific papers. Due to the multi-modal capabilities of such AI models, these new benchmarks can comprise multiple data types, potentially enhancing the applicability and transferability of these models. The continual curation of new, relevant datasets that represent the complexities of real-world chemical problems will further enhance the robustness and relevance of LLMs in chemistry.
@@
Model interpretability is a significant challenge for LLMs due to their “black-box” nature, which obscures the understanding of how predictions are made. However, innovative approaches are being developed to enhance LLMs’ interpretability. For instance, Schwaller et al. 479 and Schilter et al. 480 used information from the different multi-attention heads. While Schwaller et al. 479 connected atoms from reactants to atoms in the products, Schilter et al. 480 assigned H-NMR peaks to specific hydrogens in a molecule to indicate how spectra were comprehended, or structures deduced. Additionally, since the LLMs use language, which is intrinsically interpretable, LLMs may be incrementally modified to explain their reasoning processes directly, exemplified with tools like eXpertAI286 and or simply adjusting prompting367, 415. These methods address the critical need for transparency in the mechanism of understanding for a good prediction beyond the good prediction itself.
@@
While LLMs excel at pattern recognition, integrating explicit chemical rules and domain knowledge into these systems remains challenging. This integration is essential to make predictions that are not only statistically valid but also chemically reasonable. It was shown by Beltagy et al. 162 and Gu et al. 161 that better performance on common NLP tasks can be achieved by developing a vocabulary and pretraining on a domain-specific training corpus. While, pretrainng with domain-specific datasets that include chemical properties, reaction mechanisms, and experimental results may better capture the nuances of chemistry, the challenge to use AI to promote multi-disciplinary research remains. The Galactica LLM 155 also used special tokens for delineating chemical information, to relatively good success on chemistry tasks. Aryal et al. 481 also progress by creating an ensemble of specialist agents with different domains of knowledge, allowing them to interact to better answer the user query. Specifically, Aryal et al. 481 used agents with chemistry, physics, electrochemistry, and materials knowledge.
@@
RL has been successfully used in LLMs482, 483, 87, with a few applications also proposed for use in agents484, 485. The next frontier is applying RL to agents directly, to improve their ability on specific tasks. Bou et al. 96 provided a recent framework and example for generative molecular design when viewed as an RL problem (similar to RLHF) and some early success has been seen in applying the RLHF algorithm directly to protein language models where the reward model comes from scientific tasks.97 Neither of these are direct RL on language model agents, but are a step towards this goal.
@@
Comparing different agent systems is challenging due to the lack of robust benchmarks and evaluation schemes. Consequently, it is difficult to define what constitutes a “superhuman” digital chemist and reach a consensus on the criteria for success. This issue is similar to the ongoing discussions about defining artificial general intelligence (AGI) and the expected capabilities of cognitive architectures.486, 487 Once a reliable metric for evaluating such AI systems is established, it is crucial for the AI scientific community to set clear guidelines for conducting research. Currently, assessing success is challenging because the goals are not well defined. Building on this, we propose using Bloom’s taxonomy as a reference point for developing a metric to evaluate more complex reasoning and tool use in autonomous agents. This educational framework categorizes cognitive skills in a hierarchical manner, from basic recall to creative construction, providing a structured approach to assess higher-order thinking and reasoning capabilities in these systems. This adaptation could significantly enhance the evaluation of LLMs and autonomous agents, especially when tackling complex chemical challenges.
@@
As with all AI technologies, there are ethical considerations in deploying LLMs, including bias in model predictions and the potential misuse of AI-generated chemical knowledge. Ruan et al. 488 and Tang et al. 461 pointed out the necessity of different levels of regulation in agent development, noting that current alignment methods may be insufficient for ensuring safety, and using human evaluation to address the risks of such systems is not scalable. Therefore, the lack of specialized models for risk control and effective safety evaluations poses perhaps the greatest challenges when assuring the safety of LLMs with the ability to use tools. This underscores the urgency, and the opportunity, in automating red-teaming of agents to enhance safety protocols in AI technologies.

### v2

@@ 5.3 Agents for experiments planning
Building on the capabilities of ChemCrow and Coscientist in automating chemistry-related tasks, recent advances have focused on bridging the gap between virtual agents and physical laboratory environments. For example, Context-Aware Language Models for Science (CALMS)459, BioPlanner461, and CRISPR-GPT449 focus on giving support to researchers with wet-lab experimental design and data analysis.
@@
CALMS459 focuses on improving laboratory efficiency through the operation of instruments and management of complex experiments, employing conversational LLMs to interact with scientists during experiments. In addition, this agent can perform actions using lab equipment after lab equipment APIs have been provided to the agent as tools. CALMS was designed to enhance instrument usability and speed up scientific discovery, providing on-the-spot assistance for complex experimental setups, such as tomography scans, and enabling fully automated experiments. For instance, its capability was showcased through the operation of a real-world diffractometer. Although CALMS excelled in several tasks, a comparison between GPT-3.5 and Vicuna 1.5 revealed Vicuna’s limitations in handling tools.
@@
In contrast, BioPlanner461 significantly improves the efficiency of scientific experimentation by creating pseudocode representations of experimental procedures, showcasing AI’s capacity to streamline scientific workflows. Therefore, Rather than interacting directly with lab equipment through APIs, BioPlanner creates innovative experimental protocols that can be expanded upon within a laboratory setting. The initial step in BioPlanner’s process involves assessing the capability of LLMs to produce structured pseudocode based on detailed natural language descriptions of experimental procedures. In testing, BioPlanner successfully generated correct pseudocode for 59 out of 100 procedures using GPT-4, although the most common errors involved omitted units. Afterward, the authors used BioPlanned to generate a procedure for culturing an E.coli bacteria colony and storing it with cryopreservation, which ran successfully.
@@
Focusing on gene editing experiments, CRISPR-GPT449 is an agent developed to design experiments iteratively with constant human feedback. CRISPR-GPT449 aims to bridge the gap for non-experts by simplifying this process into manageable steps solvable by an LLM with access to useful tools. This agent operates in three modes based on user prompts: “Meta mode” provides predefined pipelines for common gene-editing scenarios; “Auto mode” uses the LLM to plan a sequence of tasks; and “Q&A mode” answers general questions about the experimental design. The authors demonstrate that based on human evaluations, CRISPR-GPT outperforms GPT-3.5 and GPT-4 in accuracy, reasoning, completeness, and conciseness. Additionally, they applied CRISPR-GPT to design real-world experiments for knocking out TGFBR1, SNAI1, BAX, and BCL2L1 in the human A375 cell line, achieving an editing efficiency of approximately 70% for each gene.
@@
Following the ideas of developing agents for automating experimental protocol generation, Ruan et al. 488 created a multi-agents system composed of 6 agents: Literature Scouter, Experiment Designer, Hardware Executor, Spectrum Analyzer, Separation Instructor, and Result Interpreter. The Large Language Models-based Reaction Development Framework (LLM-RDF)488 automates every step of the synthesis workflow. While other studies focus on the literature review489, 490, 491, HTS484, and reaction optimization492, 48, LLM-RDF can support researchers from literature search until the product purification. Using this system, the authors showed they could design a copper/TEMPO catalyzed alcohol oxidation reaction, optimize reaction conditions, engineer a scale-up, and purify the products, obtaining a yield of 86% and a purity >98% while producing 1 gram of product.
@@
Interestingly, despite acting on different fields and presenting different goals (while CALMS459 and LLM-RDF488 can autonomously use laboratory equipment, BioPlanner461 and CRISPR-GPT449 propose a protocol that needs to be carried out by a human), all studies implemented a “human-in-the-loop” approach. Maintaining the researcher as a vital piece of the development is a way to enhance reliability and mitigate agent limitations, such as errors and hallucinations. This approach also addresses risk and dual usage problems once a human can decide whether the agent’s suggestions are safe. 374, 493 On a slightly different direction, Organa370 fully automates the laboratory workload while providing feedback to the researcher and producing reports with the results, as discussed on Section 3.7.1.
@@ 5.4 Agents for automating cheminformatics tasks
Cheminformatics consists of applying information technology techniques to convert physicochemical information into knowledge. The process of solving cheminformatics problems commonly involves retrieving, processing, and analyzing chemical data.494 Getting inspiration from ChemCrow47 ideas, Chemistry Agent Connecting Tool Usage to Science (CACTUS)495 focused on assisting scientists by automating cheminformatics tasks. CACTUS automates the applications of multiple cheminformatics tools, such as property prediction and calculation, while maintaining the human-in-the-loop for molecular discovery. The authors investigated the performance of a diverse set of open-source LLMs, where Gemma-7B and Mistral-7B demonstrated superior performance against LLaMA-7B and Falcon-7B. In addition, the authors reported that adding domain-specific information in the prompt to align the agent to chemistry problems considerably increases model’s performance. For instance, predicting drug-likeness with a Gemma-7B agent improves the accuracy of $\sim$ 60% when aligning the agent. Prompt alignment improved the prediction of all properties they studied.
@@
Further illustrating the versatility of AI in scientific research and domain-specific tools usage is ChatMOF,463 which focuses on the prediction and generation of Metal-Organic Frameworks (MOFs). ChatMOF integrates MOF databases with its MOFTransformer496 predictor module, thereby showcasing the innovative use of genetic algorithms in guiding generative tasks from associated predictions. The authors showed that ChatMOF achieved an accuracy of $\sim$ 90% in search and prediction tasks while generative tasks have an accuracy of $\sim$ 70%. The genetic algorithm used by ChatMOF allows for the generation of a diverse array of MOF structures, which can be further refined based on specific properties requested by users. For instance, when prompted to, “generate structures with the largest surface area”, the system initially generated a broad distribution of structures with surface area centered in 3784 m2/g, and the GA evolves it to a narrower distribution with a peak at 5554 m2/g after only three generations. It is important to note that even though ChatMOF has access to a dataset of experimental values for MOFs, language model predictions guide their GA, and no further validation has been made. Lastly, Ansari and Moosavi 458 developed Eunomia, another domain-specific autonomous AI agent that leverages existing knowledge to answer questions about materials. Eunomia458 can use chemistry tools to access a variety of datasets, scientific papers and unstructured texts to extract and reason about material science information. The authors implemented a CoVe388 (Consistency Verification) scheme to evaluate the model’s answer and minimize hallucination. The authors showed that including CoVe increased the model’s precision by $\sim$ 20% when compared to previous methods such as an agent using ReAct only.393
@@
Promoting molecular discovery is a topic with great attention in the literature devoted to it and, as described extensively above, LLMs have leveraged a large amount of unstructured data to accelerate that discovery. Janakarajan et al. 462 discuss the advantages of using LLMs in fields such as de novo drug design, reaction chemistry, and property prediction, but they augment the LLM in IBM ChemChat, a chatbot with the capability of using common APIs and python packages commonly used daily by a cheminformatics researcher to access molecular information. ChemChat has access to tools such as Generative Toolkit for Scientific Discovery (GT4SD)497, a package with dozens of trained models generative models for science, rxn4chemistry498, a package for computing chemistry reactions tasks, HuggingMolecules499, a package developed to aggregate molecular property prediction LMs, and RDKit500, a package to manipulating molecules. Since ChemChat implements an agent in a chat-like environment, users can interactively refine design ideas. Despite being developed to target de novo drug design, ChemChat nonetheless is a multi-purpose platform that can be more broadly used for molecular discovery.
@@
In addition to the capabilities described above, LLM-based agents can empower users to tackle tasks that typically require extensive technical knowledge. In previous work, Wellawatte and Schwaller 309 and Gandhi and White 257 showed that including natural language explanations (NLE) in explainable AI (XAI) analysis can improve user understanding. More recently, Wellawatte and Schwaller 309 developed XpertAI309 to seamlessly integrate XAI techniques with LLMs to interpret and explain raw chemical data autonomously. Applying XAI techniques is usually restricted to technical experts but by integrating such techniques with an LLM-based agent to automate the workflow, the authors made XAI accessible to a wider audience.
@@
Their system receives raw data with labels for physicochemical properties. The raw data is used to compute human-interpretable descriptors and then calculate SHAP values or Z-scores for Local Interpretable Model-agnostic Explanations (LIME). By calculating SHAP values (or SHapley Addictive exPlanations) a value can be assigned to each feature, indicating its contribution to a model’s output. LIME interprets a model by making a local approximation, around a particular prediction, to indicate what factors have contributed to that prediction in the model. It may use, for example, a surrogate local linear regression fit to recognized features.257 In addition to XAI tools, XpertAI can search and leverage scientific literature to provide accessible NLEs. While ChatGPT provides scientific justifications with similar accuracy, its explanation is often too broad. On the other hand, XpertAI provides data-specific explanations and visual XAI plots to support its explanations.309 With a similar goal, Zheng et al. 501 prompted the LLM to generate explanatory rules from data.
@@
5.5 Agents for hypothesis creation
Following the agent’s classification proposed by Gao et al. 373, the studies we discussed previously lie mainly in level 1 — AI agents as a research assistant. That is, such agents can support researchers in executing predefined tasks, but they lack the autonomy to propose, test, and refine new scientific hypotheses. New research has been focusing on making agents able to refine scientists’ initial hypotheses collaboratively, which is a required skill to achieve level 2 in Gao et al. 373 classification.
the idea of an “AI scientist” who can generate new, relevant research questions (RQ) has been pursued. Wang et al. 502 developed a framework called Scientific Inspiration Machines Optimized for Novelty (SciMON). It uses LLMs to produce new scientific ideas grounded in existing literature. SciMON retrieves inspirations from past papers and iteratively refines generated ideas to optimize novelty by comparing them with prior work. Extending these ideas, Gu and Krenn 503 used LLMs to search over a knowledge graph for inspiration to propose new personalized research ideas. Aligned with this vision, Liu et al. 460 developed CoQuest, partially automating the brainstorming for new RQs process. This system uses a human-computing interface (HCI) to allow the agent to create new RQs that can be further enhanced by human feedback. They developed two strategies for RQ generation: breadth-first, where the agent generates multiple RQs simultaneously following the original user’s prompt, and depth-first, where multiple RQs are created sequentially, building on the top of the previously generated RQ. For each RQ generation, the agent implements a ReAct393 framework with tools for literature discovery, hypothesis proposition, refinement, and evaluation. Upon evaluation of 20 HCI doctoral researchers by a post-interaction survey, the breadth-first approach was preferred by 60% of the evaluators. Interestingly, despite the evaluators’ report that the breadth-first approach gave them more control and resulted in more trustworthy RQs, the depth-first had better scores for novelty and surprise. This difference might be caused by the fact that the depth-first uses its own RQ to iterate. This process can introduce new keywords that users have not considered.
Focusing on generating and testing hypotheses, ChemReasoner451 uses a domain-specific reward function and computational chemistry feedback to validate agent responses. The authors combined a Monte Carlo thought search504 for catalysis with a reward function from atomistic GNNs trained to predict adsorption energy or reaction energy barriers. While the search is responsible for exploiting literature information and allowing the model to propose new materials, the hypothetic material is further tested by the GNN. This framework was applied to suggest materials for adsorbates, biofuel catalysts, and catalysts for CO2 to methanol conversion. The LLM generated the top five catalysts for each task, with ChemReasoner significantly outperforming GPT-4 based on the reward score.
Similarly, Ma et al. 448 developed Scientific Generative Agent (SGA) to generate hypotheses and iteratively refine them through computational simulations. Initially, the LLM generates a hypothesis. In the use cases considered, it can be a code snippet or a molecule. In the sequence, a search algorithm is used to find a better initial hypothesis for solving the initial query. Finally, this hypothesis — code or molecule — is optimized using a gradient-based algorithm. Lastly, the optimization output serves as feedback to the LLM to iterate. In their molecule design task, the goal was to generate a molecule with a specified HOMO-LUMO gap. The hypothesis is a molecule, that is, a SMILES string and a set of atomic coordinates. The gap is predicted by employing UniMol505. They showed that SGA could generate molecules based on quantum mechanical properties, but the results were not validated.
@@ 6 Challenges and opportunities
Quality and availability of data are critical factors that influence the efficacy of LLMs. Indeed, scaling both the model size and the amount of training data used has proven to improve capabilities.506 However, current AI models are not trained on large amounts of chemical data, which limits their capabilities to reason about advanced chemical concepts.183
There are two types of datasets commonly used to train LLMs: unlabeled and labeled datasets. Unlabeled datasets, or pretraining data, are used during the semi-supervised training, which focuses on creating a “prior belief” about a molecule. Currently, we have huge datasets composed of hypothetical and/or theoretical data. When a model is trained on data that is not grounded in real chemical information, this might cause the model to learn a wrong prior belief.507
Labeled datasets, often used in benchmarks, also suffer from their inclusion of hypothetical and calculated data. Benchmarks are necessary for quantifying improvements in AI modeling and prediction within a competitive field. However, dominant benchmarks like MoleculeNet,57 have significant limitations that may restrict the generalizability and applicability of evolving models. In his blog, Walters 186 brings to light numerous errors and inconsistencies within the MoleculeNet data, which substantially impact model performance and reliability.184, 146, 185, 128. Walters also argues that the properties present in these benchmarks do not directly correlate with real chemistry improvement. As such, new benchmarks need to translate to practical chemistry problems directly. For instance, increasing accuracy in predicting LogP is not necessarily mapped to drugs with greater bioavailability. Some promising work has come from the Therapeutic Common Data (TDC)189, 188 includes data from actual therapeutic essays, providing a more practical foundation for model training.
The community continues to work to organize and curate datasets to make data ready for LLMs training and evaluation. Scientific benchmarks135, 456, 508, repositories with curated datasets182, and packages for model evaluation183 have been developed. However, the challenges concerning grounded truth and consistent datasets remain. With advancements in scientific document processing,339 there is now the opportunity to obtain new datasets from peer-reviewed scientific papers. Due to the multi-modal capabilities of such AI models, these new benchmarks can comprise multiple data types, potentially enhancing the applicability and transferability of these models. The continual curation of new, relevant datasets that represent the complexities of real-world chemical problems will further enhance the robustness and relevance of LLMs in chemistry.
@@
Model interpretability is a significant challenge for LLMs due to their “black-box” nature, which obscures the understanding of how predictions are made. However, innovative approaches are being developed to enhance LLMs’ interpretability. For instance, Schwaller et al. 509 and Schilter et al. 510 used information from the different multi-attention heads. While Schwaller et al. 509 connected atoms from reactants to atoms in the products, Schilter et al. 510 assigned H-NMR peaks to specific hydrogens in a molecule to indicate how spectra were comprehended, or structures deduced. Additionally, since the LLMs use language, which is intrinsically interpretable, LLMs may be incrementally modified to explain their reasoning processes directly, exemplified with tools like eXpertAI309 and or simply adjusting prompting391, 439. These methods address the critical need for transparency in the mechanism of understanding for a good prediction beyond the good prediction itself.
@@
While LLMs excel at pattern recognition, integrating explicit chemical rules and domain knowledge into these systems remains challenging. This integration is essential to make predictions that are not only statistically valid but also chemically reasonable. It was shown by Beltagy et al. 134 and Gu et al. 133 that better performance on common NLP tasks can be achieved by developing a vocabulary and pretraining on a domain-specific training corpus. While, pretrainng with domain-specific datasets that include chemical properties, reaction mechanisms, and experimental results may better capture the nuances of chemistry, the challenge to use AI to promote multi-disciplinary research remains. The Galactica LLM 126 also used special tokens for delineating chemical information, to relatively good success on chemistry tasks. Aryal et al. 511 also progress by creating an ensemble of specialist agents with different domains of knowledge, allowing them to interact to better answer the user query. Specifically, Aryal et al. 511 used agents with chemistry, physics, electrochemistry, and materials knowledge.
@@
RL has been successfully used in LLMs512, 513, 94, with a few applications also proposed for use in agents514, 515. The next frontier is applying RL to agents directly, to improve their ability on specific tasks. Bou et al. 103 provided a recent framework and example for generative molecular design when viewed as an RL problem (similar to RLHF) and some early success has been seen in applying the RLHF algorithm directly to protein language models where the reward model comes from scientific tasks.104 Neither of these are direct RL on language model agents, but are a step towards this goal.
@@
Comparing different agent systems is challenging due to the lack of robust benchmarks and evaluation schemes. Consequently, it is difficult to define what constitutes a “superhuman” digital chemist and reach a consensus on the criteria for success. This issue is similar to the ongoing discussions about defining artificial general intelligence (AGI) and the expected capabilities of cognitive architectures.516, 517 Once a reliable metric for evaluating such AI systems is established, it is crucial for the AI scientific community to set clear guidelines for conducting research. Currently, assessing success is challenging because the goals are not well defined. Building on this, we propose using Bloom’s taxonomy as a reference point for developing a metric to evaluate more complex reasoning and tool use in autonomous agents. This educational framework categorizes cognitive skills in a hierarchical manner, from basic recall to creative construction, providing a structured approach to assess higher-order thinking and reasoning capabilities in these systems. This adaptation could significantly enhance the evaluation of LLMs and autonomous agents, especially when tackling complex chemical challenges.
@@
As with all AI technologies, there are ethical considerations in deploying LLMs, including bias in model predictions and the potential misuse of AI-generated chemical knowledge. Ruan et al. 518 and Tang et al. 487 pointed out the necessity of different levels of regulation in agent development, noting that current alignment methods may be insufficient for ensuring safety, and using human evaluation to address the risks of such systems is not scalable. Therefore, the lack of specialized models for risk control and effective safety evaluations poses perhaps the greatest challenges when assuring the safety of LLMs with the ability to use tools. This underscores the urgency, and the opportunity, in automating red-teaming of agents to enhance safety protocols in AI technologies.